OpenStudy (anonymous):

A and B are events defined on a sample space, with the following probabilities. Find P(A and B). Give your answer correct to two decimal places. P(A) = 0.57 and P (B|A) = 0.38

OpenStudy (jmark):

Given P(A) = 0.57 and P (B|A) = 0.38 then P(A and B)=? we have P (B|A)=P(A and B)/P(A) 0.38=P(A and B)/0.57 P(A and B)=0.38*0.57 P(A and B)=0.2166 Learn more about Conditional Probability http://goo.gl/LPvD4k